# MLAAD: The Multi-Language Audio Anti-Spoofing Dataset

### v8

@@ Abstract
Text-to-Speech (TTS) technology offers notable benefits, such as providing a voice for individuals with speech impairments, but it also facilitates the creation of audio deepfakes and spoofing attacks. AI-based detection methods can help mitigate these risks; however, the performance of such models is inherently dependent on the quality and diversity of their training data. Presently, the available datasets are heavily skewed towards English and Chinese audio, which limits the global applicability of these anti-spoofing systems. To address this limitation, this paper presents the Multi-Language Audio Anti-Spoofing Dataset (MLAAD), version 8, created using 119 TTS models, comprising 58 different architectures, to generate 570.3 hours of synthetic voice in 40 different languages.
We train and evaluate three state-of-the-art deepfake detection models with MLAAD and observe that it demonstrates superior performance over comparable datasets like InTheWild and FakeOrReal when used as a training resource. Moreover, compared to the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. In tests across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, each excelling on four datasets.
By publishing [^1] https://deepfake-total.com/mlaad and making a trained model accessible via an interactive webserver [^2] https://deepfake-total.com/ , we aim to democratize anti-spoofing technology, making it accessible beyond the realm of specialists, and contributing to global efforts against audio spoofing and deepfakes.
@@ I Introduction
21 22 zip v2 28.02.2024 63.0 21 22 zip v3 16.04.2024 163.9 54 23 zip v4 24.09.2024 175.0 59 23 zip v5 02.11.2024 398.3 87 38 zip + Huggingface v6 26.04.2025 420.7 91 38 Huggingface v7 11.07.2025 485.3 101 40 Huggingface v8 01.10.2025 570.3 119 40 Huggingface
@@
To address these issues, we introduce and continuosly expand our MLAAD dataset (c.f. Section I), a large-scale corpus of state-of-the-art audio fakes encompassing 40 languages. This dataset, based on the M-AILABS Speech Dataset [10], includes 570.3 hours of synthesized speech, created by 119 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ III-A Synthesis Procedure
We introduce the MLAAD dataset, an extension of the M-AILABS Speech Dataset, which features audio clips in eight original languages: English, French, German, Italian, Polish, Russian, Spanish, and Ukrainian. All of these audio clips represent authentic human speech, taken from audiobooks or speeches and interviews of public figures, such as former German Chancellor Angela Merkel. We enrich this dataset with computer-generated audio clips in 40 languages.
@@ III-C Statistics
The dataset encompasses 570.3 hours of synthesized speech across 40 languages, generated using 119 state-of-the-art TTS models, comprising 58 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
Figure 2, along with Table VI and Table VII, presents the architecture, dataset, and language distribution within MLAAD, offering an overview of the TTS architectures, training datasets, and languages that constitute the dataset. Here, “training datasets” indicates the data on which the TTS systems have been trained. These include both single-speaker datasets [39, 43, 44, 62, 40], as well as larger, multi-language and multi-speaker databases [41, 42, 10]. For some of the models, the training database is either unknown, non-existent (e.g., Griffin Lim), or composed of multiple datasets. The overall composition of TTS training datasets is highly diverse and includes both professional speakers [39] and audio recorded through community-driven initiatives [42].
@@ IV-B Cross-dataset evaluation
IV-C Assessment of Synthesis Quality
To assess the quality of the synthesized audio, we utilize the automatic speaker recognition tool “Whisper” [69] to convert audio files from our dataset back into text. Then, the edit distance [70] $ED$ is computed between the original text $t$ from the M-AILABS dataset, and the text transcribed from the synthesized audio $s_{t}$:
$\text{ED}_{s}=ED(t,ASR(s_{t}))$
For a baseline, we use the original audio samples from M-AILABS $s_{o}$ to evaluate the ASR system’s accuracy. This is calculated as:
$\text{ED}_{o}=ED(t,ASR(s_{o}))$
The comparison of $\text{ED}_{s}$ with $\text{ED}_{o}$ helps to gauge the synthesized audio’s fidelity. If $\text{ED}_{s}$ is close to $\text{ED}_{o}$, it indicates that the quality of the synthesized audio is comparable to the originals.
The outcomes are detailed in Table V. In many cases, the error rates observed between the synthesized audio and the original audio are comparable. This is true especially for languages present in the original M-AILABS dataset, as well as other central European languages such as Dutch, Czech or Finish. However, in cases where the alphabet significantly diverges from the ASCII standard, such as Hindi, Maltese, Swahili or Ukrainian, we observe a noticeable increase in error rates. Despite the TTS models (and possibly the ASR systems) supporting numerous languages, a substantial disparity in quality remains evident. Prominent languages like English, Spanish or German exhibit lower error rates, whereas others encounter more challenges in accurate processing. This variation underscores the importance of enhancing support for underrepresented languages, as the current performance disparity disproportionately affects individuals from these communities.
Fake Original
Arabic 8.0 ± 5.3
Bulgarian 15.1 ± 37.8
Czech 5.8 ± 4.3
Danish 25.8 ± 13.7
Dutch 6.9 ± 5.3
German 2.4 ± 2.6 3.5 ± 7.4
Greek 16.0 ± 27.2
English 4.5 ± 4.9 2.8 ± 3.7
Estonian 7.8 ± 6.1
Finnish 2.3 ± 2.0
French 4.8 ± 5.1 4.0 ± 4.3
Hindi 18.7 ± 34.7
Hungarian 5.9 ± 4.3
Italian 9.5 ± 10.4 2.0 ± 2.0
Maltese 41.8 ± 17.4
Polish 5.8 ± 14.0 2.7 ± 3.1
Romanian 6.9 ± 5.4
Russian 4.9 ± 4.6 2.9 ± 3.6
Slovak 9.3 ± 5.9
Spanish 2.5 ± 3.1 2.1 ± 2.5
Swahili 33.2 ± 16.8
Ukrainian 23.2 ± 34.0 6.1 ± 11.3
Mean 11.9 ± 18.9 3.1 ± 5.3
TABLE V: Quality of the MLAAD v1 dataset, as measured in mean edit distance $\text{ED}_{s}$ (Fake) and $\text{ED}_{o}$ (Original), with standard deviation shown. Edit distance is computed between the original text (or translated version thereof) and the transcript of the corresponding audio file.
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 40 languages by 119 TTS models, composed of 58 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.

### v9

@@ Abstract
This paper presents the Multi-Language Audio Anti-Spoofing Dataset (MLAAD), version 9: a dataset of synthetic audio to train and evaluate audio deepfake detection models. It features 140 Text-to-Speech (TTS) models, comprising a total of 687.4 hours of synthetic voice in 51 different languages.
To evaluate this dataset, we train three state-of-the-art deepfake detection models with MLAAD and observe that it demonstrates superior performance over comparable datasets like InTheWild and FakeOrReal when used as a training resource. Moreover, compared to the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. In tests across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, each excelling on four datasets.
By publishing the dataset [^1] https://deepfake-total.com/mlaad and making a trained model accessible via an interactive webserver [^2] https://deepfake-total.com/ , we aim to democratize anti-spoofing technology, making it accessible beyond the realm of specialists, and contributing to global efforts against audio spoofing and deepfakes.
@@ I Introduction
21 22 zip v2 28.02.2024 63.0 21 22 zip v3 16.04.2024 163.9 54 23 zip v4 24.09.2024 175.0 59 23 zip v5 02.11.2024 398.3 87 38 zip + Huggingface v6 26.04.2025 420.7 91 38 Huggingface v7 11.07.2025 485.3 101 40 uggingface v8 01.10.2025 570.3 119 40 Huggingface v9 17.01.2026 687.4 140 51 Huggingface
@@
To address these issues, we introduce and continuosly expand our MLAAD dataset (c.f. Section I), a large-scale corpus of state-of-the-art audio fakes encompassing 51 languages. This dataset, based on the M-AILABS Speech Dataset [10], includes 687.4 hours of synthesized speech, created by 140 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ III-A Synthesis Procedure
We introduce the MLAAD dataset, an extension of the M-AILABS Speech Dataset, which features audio clips in eight original languages: English, French, German, Italian, Polish, Russian, Spanish, and Ukrainian. All of these audio clips represent authentic human speech, taken from audiobooks or speeches and interviews of public figures, such as former German Chancellor Angela Merkel. We enrich this dataset with computer-generated audio clips in 51 languages.
@@ III-C Statistics
The dataset encompasses 687.4 hours of synthesized speech across 51 languages, generated using 140 state-of-the-art TTS models, comprising 78 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
Figure 2, along with Table V and Table VI, presents the architecture, dataset, and language distribution within MLAAD, offering an overview of the TTS architectures, training datasets, and languages that constitute the dataset. Here, “training datasets” indicates the data on which the TTS systems have been trained. These include both single-speaker datasets [39, 43, 44, 62, 40], as well as larger, multi-language and multi-speaker databases [41, 42, 10]. For some of the models, the training database is either unknown, non-existent (e.g., Griffin Lim), or composed of multiple datasets. The overall composition of TTS training datasets is highly diverse and includes both professional speakers [39] and audio recorded through community-driven initiatives [42].
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 51 languages by 140 TTS models, composed of 78 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.